OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? twenty four, thirty, thirty five twelve, eighteen, thirty eighteen, twenty four, thirty eighteen, twenty four, thirty five

OpenStudy (anonymous):

Help please I really need this

OpenStudy (31356):

What do you think?

OpenStudy (31356):

This one requires you to use the pythagorean theorem, a^2 + b^2 = c^2 (^2 means "squared")

OpenStudy (anonymous):

i have no idea i studied but my brain is screaming HEADACHE!

OpenStudy (31356):

For the third one, 18 is a, 24 is b, and 30 is c 18^2 + 24^2 = 324 + 576 = 900 30^2 = 900 Since these are equal, these segments form a right triangle

OpenStudy (31356):

Medal? If helped.......

OpenStudy (31356):

Thanks! :D

OpenStudy (anonymous):

How to give a meadal?

OpenStudy (anonymous):

Oh i did

OpenStudy (31356):

You just did :D

OpenStudy (anonymous):

ok i figured it out

OpenStudy (31356):

:D

OpenStudy (anonymous):

How will adding the value 2 affect the mean and median of the data set 4, 4, 4, 6, 8? A. The mean and median will decrease by the same amount. B. The median will decrease more than the mean. C. The median will stay the same, but the mean will decrease. D. The mean and median will both stay the same.

OpenStudy (tennistar):

Hi @TheKING822 Post a seperate question where it says ask a question